ICM 1998, Berlin, Aug. 18–27 Abstracts of Plenary and Invited Lectures

Section:

0. Plenary Lectures

1991 MS Classification: Gallavotti, Giovanni, Universitá La Sapienza, Roma, Italy The Chaotic hypothesis and universal large deviations properties The chaotic hypothesis will be discussed and its meaning and consequences will be exemplified via some applications to Anosov systems.

ICM 1998 Berlin · Plenary and Invited Lectures